OpenStudy (anonymous):

use the quadratic formula to solve the following: x^2=-2x+6

OpenStudy (anonymous):

First re-arrange them x^2+2x-6=0 Now tell me do you know quadratic formula? :)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Okay I'll tell you then you try to solve it I'll explain you step by step :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

 Make data first this will help you alot :) The equation is x^2+2x-6=0 So in this ^a ^b ^c a=1 b=2 c=-6 Got it? :)